The world's smallest mammal is the bumblebee bat, also known as the Kitti's hog-nosed bat (or Craseonycteris thonglongyai). Such bats are roughly the size of a large bumblebee. A random sample of 25 bumblebee bats has a mean of 1.71 g and a standard deviation of 0.26 g. Assume that bumblebee bats come from the population of which weights are normally distributed. Test the claim that the mean weight of these bats is less than 1.8 g use a 1% significance level. The test statistics and the critical value are

Select one:

a. -1.73 and -2.80 respectively

b. -1.73 and 2.80 respectively

c. -1.73 and -2.49 respectively

d. None of the other answers in necessary true.

e. -1.73 and 2.49 respectively
Answer #1

Ans:

Test statistic:

t=(1.71-1.8)/(0.26/SQRT(25))

t=-1.73

df=25-1=24

critical t value=-2.49

Option c is correct.

-1.73 and -2.49 respectively